Evaluate the following prod ucts without multiplying directly:
(i) 103 x 107 (ii) 95 x 96 (iii) 104 )( 96.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) 103 x 107 • (100 + 3X100 + 7) = (100>2 + (3 + 7) x 100 + 3 )( 7 (Using identity:(.x + a)(x + b) = x2 + (a + b')x + ab] = 10000 + 1000 + 21 = 11021. 2 + (5 + 6) x 90 + 5 )( 6 [Using identity: (.x + a)(x + b) = x2 + (a + b")x + ab] ... 8100 + 990 + 30 = 9120. (iii) 104 )( 96 = (100 + 4)(100 - 4) = (100>2 -(4>2 (Using identity:(x + yXx - y) = x2 -y2] ... 10000 - 16 = 9984.